Outage: All telephony services (solved)
We are currently experiencing an outage with telephony. The outage is currently under investigation by our NOC team. Further updates will be released as soon as possible.
Please find all details below:
|Impacted services:||All telephony services|
|Affected Areas:||All areas|
Update 18/11 20:00: Full capacity is restored, multiple link resets were needed to restore all links. The problem will be investigated and corrective measures will be taken to avoid this issue in the future.
Update 18/11 14:56: The hardware replacement did not have the required effect and telephony is currently running at reduced capacity. We are still looking into solving the issue as soon as possible.
Update 18/11 13:52: The fault was found (hardware failure) and the piece has been replaced. We are following up and monitoring the issue.
We apologize for the inconveniences.
The edpnet team
*All times are listed in CET, Central European Time